find two consecutive even integers whose product is 168
Answer by mananth(16949)

find two consecutive even integers whose product is 168
lET ONE NUMBER BE X
THE OTHER NUMBER WILL BE X+2
x(x+2)= 168
x^2+2x-168=0
x^2+14x-12x-168=0
x(x+14)-12(x+14)=0
(x+14)(x-12)=0
x= 12 or -14